OpenStudy (anonymous):

-4 Times Square root of 24

OpenStudy (mathlegend):

\[-4 * \sqrt{24}\]

OpenStudy (mathlegend):

What is the largest perfect square that goes into 24? @Crea

OpenStudy (anonymous):

Help

OpenStudy (mathlegend):

Do you know your perfect squares?

OpenStudy (anonymous):

Nope

OpenStudy (mathlegend):

1 * 1 = 1 2 * 2 = 4 3 * 3 = 9 etc.... So... 1, 4, & 9 are perfect squares Do you understand?

OpenStudy (anonymous):

Can you show me the answer then I can figure it out thanks

OpenStudy (mathlegend):

Is 24 divisible by any of those perfect squares listed?

OpenStudy (mathlegend):

No, that is what I refuse to do. I will not just give answers. That is against the code of conduct. Anyone who does that gets reported for abuse.

OpenStudy (anonymous):

Sorry I used 6 x 4

OpenStudy (anonymous):

I don't believe the teacher would be looking for a decimal answer. Crea, sqrt24 could also be written as sqrt6*sqrt4 correct? What is sqrt4? 2. Thus we have 2*sqrt6. -4*2*sqrt6=-8*sqrt6.

OpenStudy (mathlegend):

Thank you.

OpenStudy (mathlegend):

So using 6 and 4 is great...

OpenStudy (mathlegend):

\[-4 * \sqrt{4*6}\] Now, the 4 is perfect... so it comes out as what @Crea ... what is the square root of 4?

OpenStudy (anonymous):

So it's-8 square root of 6

OpenStudy (mathlegend):

Yes, great.

OpenStudy (anonymous):

Thank you very much

OpenStudy (mathlegend):

You are welcome. I all about knowing you have an understanding. That is why typing out the whole process shows just about no justice. People will learn my friend. Have a great day @Crea!

OpenStudy (mathlegend):

I am all*